OpenStudy (anonymous):

Solve for the roots of x in each of the equations below. x^4 - 81 = 0 x^4 + 10x^2 + 25 = 0 x^4 - x^2 - 6 = 0

OpenStudy (anonymous):

1.) 9

OpenStudy (anonymous):

I'm stuck on the others :/

OpenStudy (anonymous):

notice that you can factor the first one as: \(\large x^4-81=(x^2+9)(x^2-9) \) so now you have the equation: \(\large (x^2+9)(x^2-9)=0 \) are you solving over the reals or complex ?

OpenStudy (anonymous):

just eliminate x^2+9 because it doesnt gv real values so x = 3,-3

OpenStudy (anonymous):

or simply x^4 = 81 so x = +3 or -3 ^^

OpenStudy (anonymous):

2nd is in the form of a^2+2ab+b^2 so its (x^2+5)^2 so so x^2+5 = 0 here you get complex number x^2 = -5

OpenStudy (anonymous):

both @ByteMe

OpenStudy (anonymous):

@ByteMe

OpenStudy (anonymous):

@minoz

OpenStudy (anonymous):

81 can be written as 9*9=3*3*3*3=3^4 x^4-3^4=0=> x=+-3

OpenStudy (anonymous):

so it wouldn't be written like (byteme) answer

OpenStudy (anonymous):

http://www.tiger-algebra.com/drill/x~4-x~2-6=0/

OpenStudy (anonymous):

the above link have step by step solution

OpenStudy (anonymous):

yaa byteme answer is correct u should further solve it

OpenStudy (anonymous):

how to further solve it? is it 3, -3?

OpenStudy (anonymous):

3 and -3 are correct there is also two more roots which are imaginary x^2+9=0=>x is +-3i

OpenStudy (anonymous):

ok thanks. Is x^2+9=0=>x is +-3i further solving it?

OpenStudy (anonymous):

yaa so ans is +3,-3,+3i,-3i

OpenStudy (anonymous):

x4 + 5x2 + 5x2 + 25=0 = (x2+5) • (x2+5)=0

OpenStudy (anonymous):

The . is the multiplication sign correct?

OpenStudy (anonymous):

just use * or (x)(y)

OpenStudy (anonymous):

ok. what do I do next on this equation?

OpenStudy (anonymous):

x^2+5=0=>x=+-sqrt(-5)=+-sqrt(5)i

OpenStudy (anonymous):

I didn't see +-sqrt(-5)=+-sqrt(5)i for the answer on the website

OpenStudy (anonymous):

My pc won't let me look at the document

OpenStudy (anonymous):

check at 2.1 in this soving a single variable equation http://www.tiger-algebra.com/drill/x~4_10x~2_25=0/

OpenStudy (anonymous):

oh I see it now

OpenStudy (anonymous):

Can you check the last one for me? (x^2 + 2) • (x^2 - 3) = 0 x^2+2 = 0 x2 = -2 x=+-sqrt-2

OpenStudy (anonymous):

@minoz

OpenStudy (anonymous):

it is +-sqrt(2)i +-sqrt(3)

OpenStudy (anonymous):

@Power2Knowledge

OpenStudy (anonymous):

Thank you